Demand Curve: P=300-Qd Supply Curve: P=30+2Qs What is the effect of a price floor at P=230?

A surplus=30

B shortage=30

C No effect

D Cannot be determined
Answer #1

Option A. Surplus=30

Explanation:

Let us put the value of P = 230 in the demand curve:

230 = 300 - Qd

or, Qd = 300 - 230 = 70

Let us put the value of P = 230 in the supply curve:

230 = 30 + 2Qs

or, 2Qs = 230 - 30

or, 2Qs = 200

or, Qs = 200/2 = 100

So, when the price is 230, quantity demanded is 70 and quantity supplied is 100.

So, there will be surplus = 100 -70 = 30